# Robustness of the Frank-Wolfe Method under Inexact Oracles and the Cost of Linear Minimization

### v1

@@ 1 Introduction
The Frank-Wolfe method is a projection-free method for constrained optimization. At each iteration, Frank-Wolfe solves a linear minimization subproblem instead of a projection, which can be advantageous on structured domains such as simplices or spectrahedra. Classical analyses guarantee an $\mathcal{O}(1/k)$ convergence rate for convex objectives Jaggi2013; FreundGrigas2013 and $\mathcal{O}(1/\sqrt{k})$ for nonconvex ones lacostejulien2016convergenceratefrankwolfenonconvex.
Despite its popularity, the robustness of Frank-Wolfe under inexact gradient information remains incompletely understood. In deterministic optimization, the $\delta$–oracle model Devolder2014 is used to bound the gradient error. We revisit this framework and derive new nonaccumulation bounds for both convex and nonconvex settings.
@@
The performance of Frank-Wolfe under inexact gradient is a very interesting problem. With unbiased gradients and bounded variance (or sub-Gaussian tails), Stochastic Frank-Wolfe variants achieve a Frank-Wolfe gap of $\mathcal{O}(\varepsilon)$ with $\mathcal{O}(1/\varepsilon^{4})$ gradient evaluations, and variance reduction accelerates finite-sum problems and can achieve the same Frank-Wolfe gap with $\mathcal{O}(1/\varepsilon^{3})$ gradient evaluations ReddiSFW2016; Goldfarb2017; LuFreund2018; Locatello2019; Zhang2020; sfyraki2025lionsmuonsoptimizationstochastic. For heavy-tailed noise, Stochastic Frank-Wolfe with clipping or robust estimation achieves high-probability guarantees Tang2022; SfyrakiWang2025.
@@
For the more common scenario, where the error does not decrease, Freund and Grigas prove an $\mathcal{O}(1/k+\delta)$ convergence of the Frank-Wolfe gapFreundGrigas2013, and we show an $\mathcal{O}(1/\sqrt{k}+\delta)$ convergence for nonconvex functions in this paper.
Sometimes we consider objective functions that are convex but non-smooth or the case when the gradients are computed at shifted points DevolderGlineurNesterov2013. Those functions may not obtain a gradient, but they can be equipped with a $(\delta,L)$ oracle Devolder2014:
@@
Since the first bound of the Frank-Wolfe gap under $(\delta,L)$- oracle, which is $\mathcal{O}(1/k+k\delta)$, has been proposedFreundGrigas2013, it has been an open problem for more than ten years whether the final guarantee of the Frank-Wolfe gap is optimal theoretically. In this paper, we improve the final guarantee of the final Frank-Wolfe gap to $\mathcal{O}(\delta)$, showing the nonaccumulation of error.
Besides robustness, the ease of computing the Linear Minimization Oracle is widely considered another major advantage of the Frank–Wolfe method, which makes it more prevalent than proximal gradient methods. However, this belief is currently supported mainly by intuition and set-specific comparisons CombettesPokutta2021; PokuttaIntro2024. Additionally, we address a recent question posed by Woodstock Woodstock2025 on whether linear minimization oracles are inherently cheaper than projection operators. We show that even coarse, approximate projections cannot outperform accurate LMOs in computational complexity.
@@ 2 Frank-Wolfe with an Inexact Oracle
Proposition 2.2 ((FreundGrigas2013, Prop. 5.1))
@@
Theorem 2.3(Nonaccumulation under $\delta$-oracle, convex caseFreundGrigas2013)
@@ Discussion.
This extends the exact-projection implication of Woodstock2025 to inexact projections: one $K$-projection at a scaled point yields an $\varepsilon$-accurate LMO. In particular, accurate linear minimization is no slower than coarse projection, uniformly over compact convex sets.

### v2

@@ 1 Introduction
The Frank-Wolfe method is a projection-free method for constrained optimization. At each iteration, Frank-Wolfe solves a linear minimization subproblem instead of a projection, which can be advantageous on structured domains such as simplices or spectrahedra. Classical analyses guarantee an $\mathcal{O}(1/k)$ convergence rate for convex objectives Jaggi (2013); Freund and Grigas (2016) and $\mathcal{O}(1/\sqrt{k})$ for nonconvex ones Lacoste-Julien (2016).
Despite its popularity, the robustness of Frank-Wolfe under inexact gradient information remains incompletely understood. In deterministic optimization, the $\delta$–oracle model Devolder et al. (2014) is used to bound the gradient error. We revisit this framework and derive new nonaccumulation bounds for both convex and nonconvex settings.
@@
The performance of Frank-Wolfe under inexact gradient is a very interesting problem. With unbiased gradients and bounded variance (or sub-Gaussian tails), Stochastic Frank-Wolfe variants achieve a Frank-Wolfe gap of $\mathcal{O}(\varepsilon)$ with $\mathcal{O}(1/\varepsilon^{4})$ gradient evaluations, and variance reduction accelerates finite-sum problems and can achieve the same Frank-Wolfe gap with $\mathcal{O}(1/\varepsilon^{3})$ gradient evaluations Reddi et al. (2016); Goldfarb et al. (2017); Lu and Freund (2018); Locatello and others (2019); Zhang and others (2020); Sfyraki and Wang (2025b). For heavy-tailed noise, Stochastic Frank-Wolfe with clipping or robust estimation achieves high-probability guarantees Tang et al. (2022); Sfyraki and Wang (2025a).
@@
For the more common scenario, where the error does not decrease, Freund and Grigas prove an $\mathcal{O}(1/k+\delta)$ convergence of the Frank-Wolfe gapFreund and Grigas (2016), and we show an $\mathcal{O}(1/\sqrt{k}+\delta)$ convergence for nonconvex functions in this paper.
Sometimes we consider objective functions that are convex but non-smooth or the case when the gradients are computed at shifted points Devolder et al. (2013). Those functions may not obtain a gradient, but they can be equipped with a $(\delta,L)$ oracle Devolder et al. (2014):
@@
Since the first bound of the Frank-Wolfe gap under $(\delta,L)$- oracle, which is $\mathcal{O}(1/k+k\delta)$, has been proposedFreund and Grigas (2016), it has been an open problem for more than ten years whether the final guarantee of the Frank-Wolfe gap is optimal theoretically.
Besides robustness, the ease of computing the Linear Minimization Oracle is widely considered another major advantage of the Frank–Wolfe method, which makes it more prevalent than proximal gradient methods. However, this belief is currently supported mainly by intuition and set-specific comparisons Combettes and Pokutta (2021); Pokutta (2024). Additionally, we address a recent question posed by Woodstock Woodstock (2025) on whether linear minimization oracles are inherently cheaper than projection operators. We show that even coarse, approximate projections cannot outperform accurate LMOs in computational complexity.
@@ 2 Frank-Wolfe with an Inexact Oracle
Proposition 2.2 ((Freund and Grigas, 2016, Prop. 5.1))
@@
Theorem 2.3(Nonaccumulation under $\delta$-oracle, convex caseFreund and Grigas (2016))
@@ Discussion.
This extends the exact-projection implication of Woodstock (2025) to inexact projections: one $K$-projection at a scaled point yields an $\varepsilon$-accurate LMO. In particular, accurate linear minimization is no slower than coarse projection, uniformly over compact convex sets.
@@ 7 Acknowledgment
8 Compliance with Ethical Standards
Conflict of interest. The authors declare that they have no competing interests.
Ethical approval. This article does not contain any studies with human participants or animals performed by any of the authors.
Informed consent. Not applicable.
9 Declarations
Funding. No funding was received for conducting this study.
Competing interests. The authors have no competing interests to declare.
Data availability. Not applicable.
Code availability. Not applicable.